OpenStudy (anonymous):

differentiate each of the following functions and simplify. 1. f(x)= (1+√x) (2+√x) 2. h(x)= (5x^2-1)^8 (1-6x^2)^4 3. h(x)= (x^2(1+x^2)^1/2)^1/2) 4. y= (1 - cosx / 1 + cosx) 5. y= tan (xsinx) 6. y= tan^-1 (a+x/1-ax) 7. y= (1-x^2) cos^-1x - x 8. y= cot^-1 (x^2-1)^1/2 + sec^-1x 9.y= x/(a^2-x^2)^1/2 - sin^-1 x/a 10. f(x)= sin^2 ( cos 2x)

OpenStudy (psymon):

Do you know how to do any of them by chance?

OpenStudy (anonymous):

1. First expand, then use the fact that: \[\frac{d}{dx}(\sqrt{x})=\frac{1}{2\sqrt{x}}\] 2. Use the chain rule, for example: \[\frac{d}{dx}(2x+3)^3=\frac{d}{dx}(2x+3)\cdot \frac{d}{du}(u^3)=2\cdot 3(2x+3)^2=6(2x+3)^2\] Where u=2x+3 3. Again, use the chain rule, but multiple times. This time, however, you will also have to use the product rule: \[\frac{d}{dx}(f(x)g(x))=f(x)\cdot\frac{d}{dx}(g(x))+g(x)\cdot\frac{d}{dx}(f(x))\] 4. You will have to use the quotient rule: \[\frac{d}{dx}(\frac{f(x)}{g(x)})=\frac{g(x)\cdot\frac{d}{dx}(f(x))-f(x)\cdot\frac{d}{dx}(g(x))}{(g(x))^2}\] 5. First remember that tan x = (sin x)/(cos x), then use product rule on x sin x, then chain rule on tan(x sin x), then quotient rule on sin(x sin x) / cos(x sin x) 6. First, note: \[\frac{d}{dx}(\tan^{-1} x)=\frac{1}{x^2 +1}\] Then use quotient rule and chain rule 7. First note: \[\frac{d}{dx}(f(x)+g(x))=\frac{d}{dx}(f(x))+\frac{d}{dx}(g(x))\] \[\frac{d}{dx}(\cot^{-1} x)=-\frac{1}{\sqrt{1-x^2}}\] And then use product rule 8. Note: \[\frac{d}{dx}(\cot^{-1} x)=-\frac{1}{x^2 +1}\] \[\frac{d}{dx}(\sec^{-1} x)=\frac{1}{x^2 \sqrt{1-\frac{1}{x^2}}}\] 9. First note: \[\frac{d}{dx}(\sin^{-1} x)=\frac{1}{\sqrt{1-x^2}}\] Then use simple chain rule for the first half 10. Use chain rule (twice), and note: \[\sin^2 x=\sin{x}\sin{x}\] So then use the product rule